OpenStudy (anonymous):

If the following system of equations was written as a matrix equation in the form AX = C,? and matrix A was expressed in the form, a=[a c] [b d] find the value of a - b + c + d. 5x+7y=7 3x-2y=9

OpenStudy (accessdenied):

So we have... 5 * x + 7 * y = 7 3 * x + (-2)*y = 9 We want to write it in a form like this: AX = C [ a c ] [ b d ] X = C By comparison, C should be the column vector of our constants on the right-hand side. X should be a column vector of our variables because multiplication of a 2x2 and 2x1 matrix results in a 2x1 matrix. A, then, will be the coefficients of the variables because of how they multiply together. [a c] [ x ] [7] [b d] [ y ] = [9] This left-side multiplies together like this: [ a(x) + c(y) ] [7 ] [ b(x) + d(y) ] = [ 9] This looks more like the original equations now. ax + cy = 7, bx + dy = 9 5x + 7y = 7 3x - 2y = 9 Can you see these comparisons? Matrix multiplication is the key piece of information here.

OpenStudy (anonymous):

Wait I got lost when you said they multiply together

OpenStudy (accessdenied):

I say, "they multiply together," referring to the matrices A and X. Is that more clear?

OpenStudy (anonymous):

mmm ok , so what would the answer be

OpenStudy (anonymous):

this is what it looks like exactly

OpenStudy (accessdenied):

ax + cy = 7, bx + dy = 9 5x + 7y = 7 3x - 2y = 9 These are the same statements. Comparing those coefficients: a=5, c=7 b=3 d=-2 These would be then used.

OpenStudy (anonymous):

oooooh ok !! I get it now:) thank you so much

OpenStudy (accessdenied):

You're welcome! :)

OpenStudy (anonymous):

i'd give you more medals if I could!

OpenStudy (anonymous):

wait because its asking for the value of all of those, do I add them all up and whatever I get should be the answer right